# Long-time soliton dynamics via a coarse-grained space-time method

### v2

@@ Abstract
We investigate the long-time dynamics of the Sine-Gordon (SG) model under a class of perturbations whose quantum field theoretic analog — via bosonization — corresponds to the massive Schwinger model describing 1+1D relativistic QED of Dirac fermions. Classical SG solutions offer critical insight into non-perturbative effects in this quantum theory, but capturing their long-time behavior poses significant numerical challenges. To address this, we extend a coarse-graining method to spacetime using a dual-mesh construction based on the Minkowski-metric. We first validate the approach against the well-studied variant of the SG model describing magnetic fluxon dynamics in Josephson transmission lines (JTLs), where analytical and numerical benchmarks exist. We then apply the method to the Schwinger-inspired SG model and uncover long-lived bound states — “Schwinger atoms” — in which a soliton is trapped by a fixed central charge. In certain regimes, the system exhibits limit cycles that give rise to positronium-like states of oppositely charged solitons, while in others such formation is suppressed. Accessing such long-time solutions requires a rigorous implementation of outgoing boundary conditions on a finite computational domain that provide radiative dissipation to allow relaxation toward states that exist only in an infinite domain. Here we provide such a construction. Our results also suggest the possibility of analog quantum simulation of relativistic quantum field theories with JTLs. These results demonstrate the utility of spatio-temporal coarse-graining methodology for probing non-perturbative structure formation in non-linear field theories.
@@ I Introduction
The Sine-Gordon (SG) equation has emerged as a fundamental model for investigating the formation and evolution of coherent electromagnetic structures in long Josephson junctions or an array of short Josephson junctions, commonly known as Josephson Transmission Lines (JTLs) [1, 2, 3, 4, 5, 6, 7, 8]. The SG model in particular describes the dynamics of supercurrent vortices, known as fluxons, which can be accelerated by bias current up to the Swihart velocity [9] — the characteristic speed at which electromagnetic waves propagate along the tunneling region. These particle-like nonlinear wave packets, along with their bound states - formed by soliton-antisoliton pairs oscillating around a common center of mass (breathers) - and their interactions - establish a bridge between the realms of condensed matter physics, quantum information science, and relativistic field theory [4, 5, 6, 7, 8, 10, 11, 12, 13, 14, 15, 16]. Beyond their well-characterized classical relativistic dynamics, it was shown that individual fluxons can exhibit intrinsically quantum behaviors [11], further adding to the rich physics that these solitons have in store. The remarkable stability and robustness of these particle-like, macroscopic quantum entities have spurred proposals for their deployment in a variety of advanced technologies. Proposed applications range from analog quantum simulation [17] and cryogenic memory [18] to single-shot detectors [19], quantum flux shuttles [20], logic circuits [21], and amplifiers [22], among many others [23, 24, 25, 26].
@@
The connection we wish to build on – between the SG model and the relativistic electrodynamics of electrons and positrons – is well-known in high-energy physics: the zero-charge sector of the massive Schwinger Model describing the quantum electrodynamics of charged fermions in 1+1D has been shown [27, 28] to be equivalent to the Quantum SG Model, but with important additional terms. This connection has played a pivotal role in exploring phenomena in quantum gauge theories that lie beyond the scope of field-theoretical perturbation methods, such as confinement [29], chiral symmetry breaking, axial anomaly [30], and fermion condensation [27]. One of these phenomena of interest here, the Sauter–Schwinger effect (also called the Schwinger effect) [31, 32], is a conjecture in Quantum Electrodynamics (QED) that suggests the possibility of spontaneous creation of electron-positron pairs out of vacuum when subjected to an electric field above a certain threshold. This prediction is significant and has been intensely investigated [33, 34, 35, 36, 37, 38] because it suggests a non-perturbative phenomenon that indicates the scale of electromagnetic field strength at which nonlinear effects in vacuum become evident. This problem has originally been studied in the context of the massive Schwinger Model [39]. Here, we are specifically interested in the long-time evolution that follows spontaneous electron-positron (e-p) creation and relaxation towards coherent, stable structures in the long-time limit of this model.
@@
A second point we wish to make in this paper is the possibility to analyze non-perturbative quantum electrodynamics with an analog simulator based on JTLs. The conditions under which such analog simulation will be interesting and relevant to phenomenology of relativistic QED are subtle, and we leave a full analysis to a future paper. In this manuscript, we focus on the Schwinger process viewed from the lens of a semiclassical analysis of the bosonized Schwinger model, and how that semiclassical model, under conditions investigated here can be realized in JTL. We are particularly interested in the formation and the electromagnetic stability of particle-like wavepackets as well as their interactions and backreaction in the presence of an “external” electric field created by an immobile and structureless charged particle at the origin (representing a proton). In the bosonized Schwinger model, the Dirac field is thought to correspond to soliton solutions of the associated sine-Gordon (SG) model [27, 28]. This correspondence has been firmly established only in the context of a closely related quantum field theory — the Thirring model [40], which crucially does not have the mass term generated by the gauge field. In this work, we explore time-dependent solutions of the exact form of the bosonized Schwinger model and in the presence of an external source, and discuss how this model can be created using a judiciously applied bias current distribution in a JTL.
@@
Radiative dissipation is another critical ingredient in modeling relaxation toward stable structures. Accurately capturing this radiative relaxation places significant demands on the size of the computational domain. An alternative technique to go around this difficulty is an open quantum system approach, where the infinite domain is divided into a system and a bath, with stringent assumptions on the bath subsystem and its coupling to the system. Here we take an alternative, third route, introducing a space-time approach to implement outgoing boundary conditions at the surface of an appropriately chosen finite computational domain. A variant of this approach has recently been shown to be compatible with the corresponding second quantized theory in the case of scalar non-relativistic electrodynamics [41, 42].
@@ II Summary of Results
We study the long-time evolution and structure formation in SG models. To do so, we first present our numerical method that builds upon DEC-QED - a coarse-graining approach that was recently developed for charged quantum fluids in multiscale heterogeneous environments [43, 44]. By incorporating the Minkowski metric into the formulation of the nonlinear wave equation, the spatial and temporal dimensions can be consistently discretized together. Combining with the structure-preserving framework of DEC-QED, this allows for an enhanced stability for long-time simulations that is also resource-efficient.
@@
To showcase the versatility of the proposed numerical tool in practical modeling of Josephson junctions, we also examine fluxon dynamics in a junction featuring narrow constrictions. In this setting, we observe signatures of Cherenkov radiation — a phenomenon previously thought to occur in long Josephson junctions only when multiple junctions are coupled together [45]. We also study dynamical creation of fluxons and breathers in a junction using external electromagnetic pulses. The efficiency is illustrated through the finding of pulse parameters that would set off the desired soliton dynamics.
To demonstrate the effectiveness of the proposed method for capturing multi-scale dynamics, we simulate the dynamical screening of a charged capacitor caused by the backreaction of e-p pairs generated via the Schwinger process [32] from vacuum. By comparing our numerical simulations with the analytical results in Ref. [34] for the massless case, we observe that the full evolution of charge and current distributions — including rapid oscillations, intermediate-time interference, and long-time asymptotic decay — is accurately reproduced. Furthermore, because the method is grounded in the rigorous framework of discrete differential forms [46], applying DEC-QED to the full spacetime grid enables precise enforcement of conservation laws, thereby improving numerical stability. In particular, the local charge-current continuity condition is exactly captured through its corresponding geometric identity, achieving machine-precision accuracy. Total energy is approximately conserved over long simulations, with numerical artifacts such as artificial micro-oscillations significantly reduced with respect to other numerical integration techniques we compare to.
@@ III Background
Figure 1: (a) 3D Schematic of a long junction in which the generalized flux variable $\varphi(x,t)$ varies along the length of the junction. (b) A view at an $xz$-slice of the junction. $\lambda_{L}$ is the London penetration depth and $\lambda_{J}$ is the Josephson penetration depth. (c) $\varphi(x)$, $\partial_{x}\varphi$ and $\partial_{t}\varphi$ of a $2\pi-$kink solution of the pure SG equation.
@@ III.1 The Sine-Gordon Equation for Fluxons
A typical (short) Josephson junction is one whose lateral dimensions are much shorter than the Josephson penetration depth ($L\!\ll\!\lambda_{J}$ in Fig. 1[47]. For a junction made of an insulator with thickness $a$ sandwiched between two superconducting electrodes having London penetration depth $\lambda_{L}$, $\lambda_{J}$ is given by $\lambda_{J}\approx\sqrt{\Phi_{0}/[2\pi\mu_{0}j_{c}\,(a+2\lambda_{L})]}$ [48], where $\Phi_{0}$ is the flux quantum, and $j_{c}$ is the critical current density. For junctions typically used in superconducting microwave circuits, $\lambda_{J}$ is in the order of a few $\mu m$, while $\lambda_{L}\approx$ few $nm$. The dynamics of Josephson junctions in an electromagnetic environment are encoded in the gauge-invariant, generalized flux variable $\varphi$ given by
@@
where the line integral is carried out from one superconducting island to the other, while $\bm{\mathcal{A}}={\bf A}-\frac{\hbar}{q}\nabla\theta$, with ${\bf A}$ being the magnetic vector potential and $\theta$ is the phase variable in the Madelung representation of the superconducting order parameter $\psi$ ($\psi=\rho e^{i\theta}$)  [49]. If one of the two lateral dimensions of the junction is sufficiently long so that the externally applied magnetic field only penetrates into a thin layer at the boundary of the junction ($L\gg\lambda_{J}$), the generalized flux variable $\varphi=\varphi(x,t)$ will vary along this long junction (see Fig. 1) and its dynamics are described by the perturbed SG equation [1, 3]
@@
In Eq. 2 above, $x$ labels the coordinate (in units of Josephson penetration depth $\lambda_{J}$) along the long dimension of the junction, and $t$ is the time variable normalized by the inverse of the Josephson plasma frequency $\omega_{p}/2\pi=c\sqrt{\mu_{0}j_{c}/\Phi_{0}}$. In Eq. 2, the perturbations to the original SG equation are characterized by the dissipative term $\alpha$ quantifying the resistive loss in the junction, and the term $\beta$ describing the amplitude of an external bias current injected into the junction which can be spatiotemporally varying.
@@
If there are no perturbations ($\alpha=\beta=0$), a fluxon corresponds to the analytical solution of Eq. 2 that describes a single soliton [1]. In general, the unperturbed SG equation permits solutions containing an arbitrary number of fluxons, antifluxons (i.e. antivortices), as well as radiation. While analytical solutions exist for the pure SG case [1] and for SG model with small perturbations [50, 51] (we discuss some analytically known solutions in Section V), the precise dynamics of solitons in the presence of significant losses and external biases can only be studied numerically.
@@ III.2 The Sine-Gordon Equation for the Bosonized Schwinger Model
The Schwinger model  [^1] Strictly speaking, the term Schwinger model traditionally refers to 1+1D QED with massless fermions, first introduced as a toy model for illustrating confinement in Ref. [54]. In this work, however, we will use the term to refer to its massive counterpart. is a 1+1D model of relativistic quantum electrodynamics of Dirac fermions coupled to the electromagnetic field. The Lagrangian is given by
@@
where $\psi$ is the fermion spinor field, $m$ is the mass of the fermion, $D_{\mu}=\partial_{\mu}-ieA_{\mu}$ is the gauge covariant derivative, and $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the electromagnetic field strength tensor, which in 1+1D has only one independent component, $F^{\mu\nu}=\epsilon^{\mu\nu}F^{01}$ (where $\epsilon^{\mu\nu}$ is the 1+1D Levi-Civita tensor, so that $\epsilon^{01}=1$), identified as the electric field $F^{01}=-F^{10}=E(x,t)$. The Schwinger model can be equivalently rewritten as a theory of bosons where components of the Dirac field $\psi$ are mapped to exponentials of the corresponding boson fields, i.e. the vertex operators in the bosonic theory [27, 28, 53]. In particular, the bosonized equivalence of the Schwinger model is described by the following Lagrangian
@@
where $\varphi(x,t)$ is a Hermitian scalar field satisfying bosonic commutation relations, $g=\frac{e}{\sqrt{\pi}}$, and $\kappa=\frac{m}{\pi\epsilon}$ where $\epsilon$ is a UV cutoff scale in energy. The term proportional to $\kappa$ originates from the fermionic mass term in Eq. (3), while the term proportional to $g$ — as will become clear once the gauge field is integrated out (see Eq. (7)) — leads to a dynamically generated mass for the scalar field $\varphi$. Intuitively, this can be traced back to earlier analyses [54, 55, 27] of the exactly solvable massless fermion case, which revealed that the fundamental excitations are not isolated charged particles but rather neutral, dipole-like mesonic states. These mesons acquire mass through their hybridization with the gauge field. A concise summary of the relevant bosonization framework is provided in Appendix B.
@@
$\partial_{\mu}F^{\mu\nu}=g\epsilon^{\mu\nu}\partial_{\mu}\varphi=j_{\phi}^{\nu}$ (6)
where $j_{\phi}^{\nu}$ represents the conserved current in the bosonic picture. From this, Eq. 6 can be integrated to show that $E=g\varphi$ and plugged into Eq. 5 to get a single equation of motion for the system:
$(\partial_{t}^{2}-\partial_{x}^{2})\varphi+g^{2}\phi+\sqrt{\pi}\kappa\sin(2\sqrt{\pi}\varphi)=0$ (7)
@@
[…]
Therefore, the boundary conditions on $\varphi$ restrict us to a subset of the state space in which the total fermion number (fermions minus antifermions) is fixed. The system evolves in time by creating and annihilating electron-positron pairs. An external electric field applied to a system in the vacuum state creates electron-positron pairs through the Schwinger process [56], and the electric fields between the pairs will have a backreaction effect on the original external field [34]. In Sections VI.1, VI.2.1, we consider situations where the electric field is created by a central point source   [^2] The original Schwinger process was proposed as a non-perturbative effect in 3+1D QED. Throughout this work, references to the Schwinger process and to the creation or annihilation of electron-positron pairs should be understood as their analogous counterparts in the context of the 1+1D model..
@@
Key insight into the non-perturbative aspects of quantum field theories of the type in Eq. (4) can be gained through a semiclassical approach [58, 59] which involves analyzing solutions of the corresponding classical field theory. In the remainder of this paper, we carry out such an analysis in detail.
In addition to having the same pure SG terms, Eqs. 2 and 5 differ in the additional perturbing terms, which greatly affect how the systems evolve over time and their steady states. Nevertheless, as we will see, the solutions of Eq. (5) still feature solitary waves. In the bosonized picture of the Schwinger model, the corresponding kink solution in $\varphi$ represents a transition between two discrete vacua, which we will constrain to be equal to a single electron charge in the case of the central charge problem we analyze in Section VI.2.1. Therefore, a SG soliton here will be a finite-size field packet that carries exactly one unit of charge [40].

### v3

@@ Abstract
We investigate the long-time dynamics of the Sine-Gordon (SG) model under a class of perturbations whose quantum field theoretic analog – via bosonization – corresponds to the massive Schwinger model describing 1+1D relativistic QED of Dirac fermions. Classical SG solutions offer critical insight into non-perturbative effects in this quantum theory, but capturing their long-time behavior poses significant numerical challenges. To address this, we extend a coarse-graining method to spacetime using a dual-mesh construction based on the Minkowski-metric. We first validate the approach against the well-studied variant of the SG model describing magnetic fluxon dynamics in Josephson transmission lines (JTLs), where analytical and numerical benchmarks exist. We then apply the method to the Schwinger-inspired SG model and uncover long-lived bound states – “Schwinger atoms” – in which a soliton is trapped by a fixed central charge. In certain regimes, the system exhibits limit cycles that give rise to positronium-like states of oppositely charged solitons, while in others such formation is suppressed. Accessing such long-time solutions requires a rigorous implementation of outgoing boundary conditions on a finite computational domain that provide radiative dissipation to allow relaxation toward states that exist only in an infinite domain. Here we provide such a construction. Our results also suggest the possibility of analog quantum simulation of relativistic quantum field theories with JTLs. These results demonstrate the utility of spatio-temporal coarse-graining methodology for probing non-perturbative structure formation in non-linear field theories.
@@ I Introduction
The Sine-Gordon (SG) equation has emerged as a fundamental model for investigating the formation and evolution of coherent electromagnetic structures in long Josephson junctions or an array of short Josephson junctions, commonly known as Josephson Transmission Lines (JTLs) [mclaughlin1978_SG, scott1976JJSG, ustinov1998solitons, berdiyorov2014parametric, fujimori2016field, sheikhzada2019instability, stern2019fractional, krasnov2020josephson]. The SG model in particular describes the dynamics of supercurrent vortices, known as fluxons, which can be accelerated by bias current up to the Swihart velocity [swihart1961] — the characteristic speed at which electromagnetic waves propagate along the tunneling region. These particle-like nonlinear wave packets, along with their bound states - formed by soliton-antisoliton pairs oscillating around a common center of mass (breathers) - and their interactions - establish a bridge between the realms of condensed matter physics, quantum information science, and relativistic field theory [berdiyorov2014parametric, fujimori2016field, sheikhzada2019instability, stern2019fractional, krasnov2020josephson, ustinov1992dynamics, wallraff2003singlevortex, fedorov2013josephson, sato2022quasiparticle, wildermuth2022fluxons, laub1995lorentz, gulevich2008timedilation]. Beyond their well-characterized classical relativistic dynamics, it was shown that individual fluxons can exhibit intrinsically quantum behaviors [wallraff2003singlevortex], further adding to the rich physics that these solitons have in store. The remarkable stability and robustness of these particle-like, macroscopic quantum entities have spurred proposals for their deployment in a variety of advanced technologies. Proposed applications range from analog quantum simulation [petrescu2018qsim] and cryogenic memory [kalashnikov2024memory] to single-shot detectors [soloviev2014detector], quantum flux shuttles [fulton1973fluxshuttle], logic circuits [nakajima1976logic_circuit], and amplifiers [nordman1995amplifier], among many others [devoret2013sc_review, welp2013emitters_review, golod2021phaseshifter, ghosh2024diode].
@@
The connection we wish to build on – between the SG model and the relativistic electrodynamics of electrons and positrons – is well-known in high-energy physics: the zero-charge sector of the massive Schwinger Model describing the quantum electrodynamics of charged fermions in 1+1D has been shown [coleman1975, coleman1976] to be equivalent to the Quantum SG Model, but with important additional terms. This connection has played a pivotal role in exploring phenomena in quantum gauge theories that lie beyond the scope of field-theoretical perturbation methods, such as confinement [wilson1974confinement], chiral symmetry breaking, axial anomaly [manton1985schwinger], and fermion condensation [coleman1975]. One of these phenomena of interest here, the Sauter–Schwinger effect (also called the Schwinger effect) [sauter1931, Schwinger1951], is a conjecture in Quantum Electrodynamics (QED) that suggests the possibility of spontaneous creation of electron-positron pairs out of vacuum when subjected to an electric field above a certain threshold. This prediction is significant and has been intensely investigated [schutzhold2008dynamically, chu2010capacitordischarge, semenoff2011holographic, sonner2013holographic, otto2015dynamical, otto2017afterglow] because it suggests a non-perturbative phenomenon that indicates the scale of electromagnetic field strength at which nonlinear effects in vacuum become evident. This problem has originally been studied in the context of the massive Schwinger Model [schwinger1962model]. Here, we are specifically interested in the long-time evolution that follows spontaneous electron-positron (e-p) creation and relaxation towards coherent, stable structures in the long-time limit of this model.
@@
A second point we wish to make in this paper is the possibility to analyze non-perturbative quantum electrodynamics with an analog simulator based on JTLs. The conditions under which such analog simulation will be interesting and relevant to phenomenology of relativistic QED are subtle, and we leave a full analysis to a future paper. In this manuscript, we focus on the Schwinger process viewed from the lens of a semiclassical analysis of the bosonized Schwinger model, and how that semiclassical model, under conditions investigated here can be realized in JTL. We are particularly interested in the formation and the electromagnetic stability of particle-like wavepackets as well as their interactions and backreaction in the presence of an “external” electric field created by an immobile and structureless charged particle at the origin (representing a proton). In the bosonized Schwinger model, the Dirac field is thought to correspond to soliton solutions of the associated sine-Gordon (SG) model [coleman1975, coleman1976]. This correspondence has been firmly established only in the context of a closely related quantum field theory — the Thirring model [colemanthirring], which crucially does not have the mass term generated by the gauge field. In this work, we explore time-dependent solutions of the exact form of the bosonized Schwinger model and in the presence of an external source, and discuss how this model can be created using a judiciously applied bias current distribution in a JTL.
@@
Radiative dissipation is another critical ingredient in modeling relaxation toward stable structures. Accurately capturing this radiative relaxation places significant demands on the size of the computational domain. An alternative technique to go around this difficulty is an open quantum system approach, where the infinite domain is divided into a system and a bath, with stringent assumptions on the bath subsystem and its coupling to the system. Here we take an alternative, third route, introducing a space-time approach to implement outgoing boundary conditions at the surface of an appropriately chosen finite computational domain. A variant of this approach has recently been shown to be compatible with the corresponding second quantized theory in the case of scalar non-relativistic electrodynamics [malekakhlagh2016non, kanupaper].
@@ II Summary of Results
We study the long-time evolution and structure formation in SG models. To do so, we first present our numerical method that builds upon DEC-QED - a coarse-graining approach that was recently developed for charged quantum fluids in multiscale heterogeneous environments [dec-qed, pham2023spectral]. By incorporating the Minkowski metric into the formulation of the nonlinear wave equation, the spatial and temporal dimensions can be consistently discretized together. Combining with the structure-preserving framework of DEC-QED, this allows for an enhanced stability for long-time simulations that is also resource-efficient.
@@
To showcase the versatility of the proposed numerical tool in practical modeling of Josephson junctions, we also examine fluxon dynamics in a junction featuring narrow constrictions. In this setting, we observe signatures of Cherenkov radiation — a phenomenon previously thought to occur in long Josephson junctions only when multiple junctions are coupled together [goldobin1998cherenkov]. We also study dynamical creation of fluxons and breathers in a junction using external electromagnetic pulses. The efficiency is illustrated through the finding of pulse parameters that would set off the desired soliton dynamics.
To demonstrate the effectiveness of the proposed method for capturing multi-scale dynamics, we simulate the dynamical screening of a charged capacitor caused by the backreaction of e-p pairs generated via the Schwinger process [Schwinger1951] from vacuum. By comparing our numerical simulations with the analytical results in Ref. [chu2010capacitordischarge] for the massless case, we observe that the full evolution of charge and current distributions — including rapid oscillations, intermediate-time interference, and long-time asymptotic decay — is accurately reproduced. Furthermore, because the method is grounded in the rigorous framework of discrete differential forms [hirani2003discrete], applying DEC-QED to the full spacetime grid enables precise enforcement of conservation laws, thereby improving numerical stability. In particular, the local charge-current continuity condition is exactly captured through its corresponding geometric identity, achieving machine-precision accuracy. Total energy is approximately conserved over long simulations, with numerical artifacts such as artificial micro-oscillations significantly reduced with respect to other numerical integration techniques we compare to.
@@ III Background
Figure 1: (a) 3D Schematic of a long junction in which the generalized flux variable $\varphi(x,t)$ varies along the length of the junction. (b) A view at an xzxz-slice of the junction. $\lambda_{L}$ is the London penetration depth and $\lambda_{J}$ is the Josephson penetration depth. (c) $\varphi(x)$, $\partial_{x}\varphi$ and $\partial_{t}\varphi$ of a $2\pi$-kink solution of the pure SG equation.
@@ III.1 The Sine-Gordon Equation for Fluxons
In this section, we provide some background information on how the Sine-Gordon equation arises in long Josephson junctions. A more simple model, the short Josephson junction, is one whose lateral dimensions are much shorter than the Josephson penetration depth ($L\!\ll\!\lambda_{J}$ in Fig. 1) [buckel2008superconductivity]. For a junction made of an insulator with thickness $a$ sandwiched between two superconducting electrodes having London penetration depth $\lambda_{L}$, $\lambda_{J}$ is given by $\lambda_{J}\approx\sqrt{\Phi_{0}/[2\pi\mu_{0}j_{c}\,(a+2\lambda_{L})]}$ [weihnacht1969influence], where $\Phi_{0}$ is the flux quantum, and $j_{c}$ is the critical current density. For junctions typically used in superconducting microwave circuits, $\lambda_{J}$ is in the order of a few $\mu m$, while $\lambda_{L}\approx$ few $nm$. The dynamics of Josephson junctions in an electromagnetic environment are encoded in the gauge-invariant, generalized flux variable $\varphi$ given by
@@
where the line integral is carried out from one superconducting island to the other, while $\bm{\mathcal{A}}={\bf A}-\frac{\hbar}{q}\nabla\theta$, with $\bf A$} being the magnetic vector potential and $\theta$ is the phase variable in the Madelung representation of the superconducting order parameter $\psi$ ($\psi=\rho e^{i\theta}$) [madelung]. In this case, the generalized flux variable $\varphi(t)$ will be described by
$\partial_{t}^{2}\varphi+\alpha\partial_{t}\varphi+\sin\varphi=-\beta$ (2)
where the dissipative term $\alpha$ quantifies the resistive loss in the junction, and the term $\beta$ describes the amplitude of an external bias current injected the junction.
If one of the two lateral dimensions of the junction is sufficiently long such that the externally applied magnetic field only penetrates into a thin layer at the boundary of the junction ($L\gg\lambda_{J}$), it is called a long Josephson junction and the generalized flux variable $\varphi=\varphi(x,t)$ will vary in $x$ along the junction (see Fig. 1) and its dynamics are described by the perturbed SG equation [mclaughlin1978_SG, ustinov1998solitons]
@@
In Eq. 3 above, $x$ labels the coordinate (in units of Josephson penetration depth $\lambda_{J}$) along the long dimension of the junction, and $t$ is the time variable normalized by the inverse of the Josephson plasma frequency $\omega_{p}/2\pi=c\sqrt{\mu_{0}j_{c}/\Phi_{0}}$. For the remainder of the text, we will specifically focus on the long junction. For a long junction with overlap geometry, $\beta$ models the bias current density across its length and can be spatiotemporally varying. For the inline configuration, the bias is injected at the junction endpoints and enters the model via boundary conditions on the sine-Gordon equation [parmentier1993solitons].
@@
If there are no perturbations ($\alpha=\beta=0$), a fluxon corresponds to the analytical solution of Eq. 3 that describes a single soliton [mclaughlin1978_SG]. In general, the unperturbed SG equation permits solutions containing an arbitrary number of fluxons, antifluxons (i.e. antivortices), as well as radiation. While analytical solutions exist for the pure SG case [mclaughlin1978_SG] and for SG model with small perturbations [rubinstein1970SG, fogel1977perturbedSG] (we discuss some analytically known solutions in Section V), the precise dynamics of solitons in the presence of significant losses and external biases can only be studied numerically. Note that although throughout this paper we refer to any additional terms added on to the Sine-Gordon model as “perturbations,” we do not use any perturbation theory to study the systems discussed in this text.
@@ III.2 The Sine-Gordon Equation for the Bosonized Schwinger Model
The Schwinger model  [^1] Strictly speaking, the term Schwinger model traditionally refers to 1+1D QED with massless fermions, first introduced as a toy model for illustrating confinement in Ref. [schwinger1962]. In this work, however, we will use the term to refer to its massive counterpart. is a 1+1D model of relativistic quantum electrodynamics of Dirac fermions coupled to the electromagnetic field. The Lagrangian is given by
@@
where $\psi$ is the fermion spinor field, $m$ is the mass of the fermion, $D_{\mu}=\partial_{\mu}-ieA_{\mu}$ is the gauge covariant derivative, and $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the electromagnetic field strength tensor, which in 1+1D has only one independent component, $F^{\mu\nu}=\epsilon^{\mu\nu}F^{01}$ (where $\epsilon^{\mu\nu}$ is the 1+1D Levi-Civita tensor, so that $\epsilon^{01}=1$), identified as the electric field $F^{01}=-F^{10}=E(x,t)$. The Schwinger model can be equivalently rewritten as a theory of bosons where components of the Dirac field $\psi$ are mapped to exponentials of the corresponding boson fields, i.e. the vertex operators in the bosonic theory [coleman1975, coleman1976, mandelstam1975]. In particular, the bosonized equivalence of the Schwinger model is described by the following Lagrangian
@@
where $\varphi(x,t)$ is a Hermitian scalar field satisfying bosonic commutation relations, $g=\frac{e}{\sqrt{\pi}}$, and $\kappa=\frac{m}{\pi\epsilon}$ where $\epsilon$ is a UV cutoff scale in energy. The term proportional to $\kappa$ originates from the fermionic mass term in Eq. (4), while the term proportional to $g$ — as will become clear once the gauge field is integrated out (see Eq. (8)) — leads to a dynamically generated mass for the scalar field $\varphi$. Intuitively, this can be traced back to earlier analyses [schwinger1962, lowenstein1971, coleman1975] of the exactly solvable massless fermion case, which revealed that the fundamental excitations are not isolated charged particles but rather neutral, dipole-like mesonic states. These mesons acquire mass through their hybridization with the gauge field. A concise summary of the relevant bosonization framework is provided in Appendix B.
@@
$\partial_{\mu}F^{\mu\nu}=g\epsilon^{\mu\nu}\partial_{\mu}\varphi=j_{\varphi}^{\nu}$ (7)
where $j_{\varphi}^{\nu}$ represents the conserved current in the bosonic picture. From this, Eq. 7 can be integrated to show that $E=g\varphi$ and plugged into Eq. 6 to get a single equation of motion for the system:
$(\partial_{t}^{2}-\partial_{x}^{2})\varphi+g^{2}\varphi+\sqrt{\pi}\kappa\sin(2\sqrt{\pi}\varphi)=0$ (8)
@@
$Q=-g\int_{-L}^{L}dx\,\partial_{x}\varphi=-g\Big[\varphi(L)-\varphi(-L)\Big]$ (9)
Therefore, the boundary conditions on $\varphi$ restrict us to a subset of the state space in which the total fermion number (fermions minus antifermions) is fixed. The system evolves in time by creating and annihilating electron-positron pairs. An external electric field applied to a system in the vacuum state creates electron-positron pairs through the Schwinger process [Schwinger1951], and the electric fields between the pairs will have a backreaction effect on the original external field [chu2010capacitordischarge]. In Sections VI.1, VI.2.1, we consider situations where the electric field is created by a central point source  [^2] The original Schwinger process was proposed as a non-perturbative effect in 3+1D QED. Throughout this work, references to the Schwinger process and to the creation or annihilation of electron-positron pairs should be understood as their analogous counterparts in the context of the 1+1D model..
@@
Key insight into the non-perturbative aspects of quantum field theories of the type in Eq. (5) can be gained through a semiclassical approach [dashen1974, dashen1975] which involves analyzing solutions of the corresponding classical field theory. In the remainder of this paper, we carry out such an analysis in detail.
In addition to having the same pure SG terms, Eqs. 3 and 6 differ in the additional perturbing terms, which greatly affect how the systems evolve over time and their steady states. Nevertheless, as we will see, the solutions of Eq. (6) still feature solitary waves. In the bosonized picture of the Schwinger model, the corresponding kink solution in $\varphi$ represents a transition between two discrete vacua, which we will constrain to be equal to a single electron charge in the case of the central charge problem we analyze in Section VI.2.1. Therefore, a SG soliton here will be a finite-size field packet that carries exactly one unit of charge [colemanthirring].